OpenStudy (anonymous):

A phone company offers two monthly plans. Plan A costs $16 plus an additional $0.10 for each minute of calls. Plan B costs $12 plus an additional $0.15 for each minute of calls. For what amount of calling do the two plans cost the same? What is the cost when the two plans cost the same? O.o....word math problems are just bleh...help please?

OpenStudy (anonymous):

Let the amount of calling be x. 16 + 0.10x = 12 + 0.15x Solve this linear equation and you will get the value of x. Substitute the value of x in any one of the equations and you will get the amount when the two plans cost the same. Just have a try!!

OpenStudy (anonymous):

Sure :) will you check my answer when I try it out, please?

OpenStudy (anonymous):

sure!!

OpenStudy (anonymous):

So I got x = 4/5...

OpenStudy (anonymous):

so far so good...bad?

OpenStudy (anonymous):

I over think things a lot...fair warning

OpenStudy (anonymous):

See, 0.05x = 4 x = 400/5 = 80

OpenStudy (anonymous):

OH I think I got it! so the amount of minutes is 80 and the cost is $24!!!

OpenStudy (anonymous):

Yes!!! You got it!! Congo!!

OpenStudy (anonymous):

congo? I like that! what does it mean?

OpenStudy (anonymous):

Once I kind of have the equation set up it gets a bit easier for me...but I hate word problems...math in general but seriously...word problems and fractions -.-

OpenStudy (anonymous):

congratulations!!

OpenStudy (anonymous):

itzz a bit difficult to transform word problems into equations, but you gotta practice them too.

OpenStudy (anonymous):

Could I give you another one for you to set up for me to solve and then check my answer? I'm doing some practice problems right now trying to learn this stupid thing...

OpenStudy (anonymous):

Yeah..!!

OpenStudy (anonymous):

Jose is going to rent a truck for one day. There are two companies he can choose from, and they have the following prices. Company A charges $90 and allows unlimited mileage. Company B has an initial fee of $55 and charges an additional $0.70 for every mile driven. For what mileages will Company A charge less than Company B? Use m for the number of miles driven, and solve your inequality for m

OpenStudy (anonymous):

See , Given that 90 should be less than 55 + 0.7m So, 90 < 55 + 0.7m Just solve this inequality.

OpenStudy (anonymous):

m = 50?

OpenStudy (anonymous):

See . It is a inequality. You cant get a equality sign. It is m > 50 When m>50 A charges less than B

OpenStudy (anonymous):

is that really it? god I wish online school explained it in an easier way...

OpenStudy (anonymous):

can we try another? I'm almost done and I'm actually really LEARNING this :D

OpenStudy (anonymous):

I promise to write you a hell of a testimonial :3

OpenStudy (anonymous):

ohkk fi9

OpenStudy (anonymous):

fi9?

OpenStudy (anonymous):

fine

OpenStudy (anonymous):

haha I fail at understanding the lingo of the internet...that's what I get for being an english nerd...it's also why my math is abysmal

OpenStudy (anonymous):

Hong is putting money into a checking account. Let y represent the total amount of money in the account (in dollars). Let x represent the number of weeks Hong has been adding money. Suppose that x and y are related by the equation 40x + 350 = y . Answer the questions below. Note that a change can be an increase or a decrease. For an increase, use a positive number. For a decrease, use a negative number. What is the change per week in the amount of money in the account? What was the starting amount of money in the account?

OpenStudy (anonymous):

For the change per week , we calculate it by taking 1st and 2nd week amounts,i.e, substitute x =1 and x=2 in the given equation. Then subtract the values. Just try it. And tell me the answer after you are sure you understood it.

OpenStudy (anonymous):

x = 1 and x = 2 or did you mean y? we substitute x twice?

OpenStudy (anonymous):

You get a value of y right. Let the value of y in the first week be y1 and in the second week be y2. You have been asked to find the change per week i.e, Amount in second week - Amount in first week = y2 - y1

OpenStudy (anonymous):

You got it??

OpenStudy (anonymous):

I think I got y = 390 and y = 430 and then subtracted that gave me y = 40

OpenStudy (anonymous):

You are right. Did you understand how it came?

OpenStudy (anonymous):

Yep, you explained it well :)

OpenStudy (anonymous):

Now for the second question, you have to find the money at the starting ,i.e, when the number of weeks is 0. So tell me what you would do?

OpenStudy (anonymous):

substitute 0 for x?

OpenStudy (anonymous):

or is it 1?

OpenStudy (anonymous):

hellooooo over thinking....

OpenStudy (anonymous):

Yes!!! you got it. Just substitute x = 0 in the equation and you will get the answer

OpenStudy (anonymous):

OK then yay I did get it the first time!

OpenStudy (anonymous):

so 350 is the answer?

OpenStudy (anonymous):

Yup..!!

OpenStudy (anonymous):

BAM I'm learning this dumb stuff that I don't like but is probably still very important!

OpenStudy (anonymous):

It surely is very important!!

OpenStudy (anonymous):

that's it for the word problems! thank you so much! I;m off to take my test :D

OpenStudy (anonymous):

All the best!!!